OpenStudy (minthia):

the lateral area of a cone is 555pi cm^2 the radius is 15 cm what is the slant height?

563blackghost (563blackghost):

To find the Lateral Area of a cone you use...\[LA = \pi r l\] since we know the lateral area and the radius we input these numbers into the formula....\[555picm^2=\pi 15 l\] what do we do to find the slant height?...

OpenStudy (minthia):

im not sure wouldn't I multiply if that's wrong PLEASE don't laugh. I struggle so much.

563blackghost (563blackghost):

We will first simplify 555pi and pi x 15 x l... \[555 \times \pi = (\pi \times 15) \times l\]

563blackghost (563blackghost):

So what is 555 times pi?

563blackghost (563blackghost):

@minthia

OpenStudy (minthia):

ok I got 174470?

OpenStudy (minthia):

4710?

563blackghost (563blackghost):

Not quite 555 times pi is equal to 1742.7.... We just simplified LA total, now we need to simplify pi by 15....\[1742.7 = ( 3.14 \times 15) l\]

OpenStudy (minthia):

47.10 do I subtract that from 1742.7

563blackghost (563blackghost):

Thats Correct...31.4 times 15 is 47.1....you will need to divide 1742.7 by 47.1....\[1742.7 \div 47.1=\]

563blackghost (563blackghost):

3.14*

OpenStudy (minthia):

37?

563blackghost (563blackghost):

Correct! it is actually 36.98 but rounded it off and it will be 37 xD ..... Great Job! So the slant height is 37 rounded or 36.98 exact....

OpenStudy (minthia):

you are AMAZING things my teacher couldn't explain and you did and I understood it. thank you sooooo much

563blackghost (563blackghost):

Np xD I am glad I was helpful ! Please press "Best Response" If I was helpful to you...

563blackghost (563blackghost):

O you already did xD Thank You!

OpenStudy (minthia):

you bet